OpenStudy (anonymous):

ok... here's a quadratic equation example... could someone do it and show there work, because no matter how people explain it... I DON"T GET IT.... So here it is... x^2-2x-24=0

OpenStudy (amistre64):

the name quadratic, comes from the old days when people used words to describe certain events; in this case the word for "we squared it and got this" was quadratic

OpenStudy (amistre64):

4^2 is a quadric (x-2)^2 is a quadratic

OpenStudy (amistre64):

quad means four; and a square has 4 sides .... that sort of logic

OpenStudy (amistre64):

what this means to me is; that a quadratic equation can be broken down into 2 baser parts; called factors. and when those factors are multiplied together, we get a product that returns us to the original form

OpenStudy (amistre64):

thats the basics as i see them, now how do we apply that to your question?

OpenStudy (amistre64):

x^2 -2x -24 = 0 ; thjis tells us that something times something else = 0 for starters

OpenStudy (amistre64):

we need to factor, or break the quadratic into its baser parts; turn it into multiplication so that we can determine its behaviour more easily

OpenStudy (amistre64):

if a*b = 0; then either a=0 or b=0 right?

OpenStudy (amistre64):

a(0) = 0 or 0(b) = 0 or 0*0=0 right?

OpenStudy (anonymous):

i think... :/

OpenStudy (amistre64):

these are basic concepts that you prolly already know; so what is it specifically that you need help on in order to factor: x^2-2x-24 = 0

OpenStudy (anonymous):

i am JUST learning all this in my algebra class which i am almost finished with. The lessons are online and I am having a hard time following it. So I pretty much need to know what steps are involved in solving it. Like what you actually do one your scratch paper, that kind of thing. Thanks!

OpenStudy (amistre64):

if it factors nice and pretty like; i mean it doesnt contain square roots and other terrible math notations; the process i use goes like this ... you ready?

OpenStudy (amistre64):

step1; factor any common factors out of it; this has none so we can move to the next step

OpenStudy (amistre64):

step 2; notice the last number: x^2-2x -24 ^^^ this one it is a negative number which tell me im gonna need to do some subtraction

OpenStudy (anonymous):

what you actually write on your paper, assuming these factor, is \[(x + ?)(x + ??)=0\] and fill the ? spots with two numbers whose product is -24 and whose sum is -2

OpenStudy (amistre64):

step 3; notice the middle term: x^2 -2x -24 ^^^ its negative which tells me the larger factor is negative, more on that soon

OpenStudy (amistre64):

step 4; find factors of 24 that subtract to get 2 1,24; 24 - 1 = 23 ... not it 2,12; 12 - 2 = 10 .... not it 4,6 ; 6 - 4 = 2 ..... thats the one

OpenStudy (amistre64):

step 5; set up your factors: (x 6)(x 4) the larger one is negative as we clued in earlier so... (x - 6) (x + 4) are our factors

OpenStudy (amistre64):

now tell me where you get lost in this :)

OpenStudy (anonymous):

Okay, i think I get it... Could I try another one and you could tell me if i get it right? That would rock.

OpenStudy (amistre64):

i could, but we are one or two steps away from finishing up this one

OpenStudy (anonymous):

oh ok sorrry

OpenStudy (amistre64):

once we got factored forms: we solve for our zeroes (x - 6) (x + 4) = 0 , recall above i stated, that it equals zero when one of the other part is 0 (0) (x+4) = 0 (x-6) (0) = 0 so..... when does: x-6 = 0 and when does: x+4 = 0

OpenStudy (amistre64):

we get 2 answers for our solution; x = 6 or x=-4 will make our equation equal zero

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

yes... :)

OpenStudy (amistre64):

good, whats your next question then :)

OpenStudy (anonymous):

0 = 6x2 − 10x − 4

OpenStudy (amistre64):

same steps; with a twist ....

OpenStudy (amistre64):

step 1 is factor out commons; i see at least a 2 that can be factored out, do you?

OpenStudy (amistre64):

0 = 2( 3x^2 −5x −2) right?

OpenStudy (anonymous):

yes...

OpenStudy (amistre64):

step 2; notice the sign for the end, that -2 part; tells us we are going to subtract out factors step 3; notice the middle term is a -5, the negative tells us the larger factor is going to be negative

OpenStudy (amistre64):

step 4; this is the twist, since we got something other than a "1" in front, we multiply it to the end to get something to factor... 3(2) = 6 ; now sort thru the factors of 6 that "subtract" to get 5. 1,6; 6 - 1 = 5 ... thats was nice, first try

OpenStudy (amistre64):

step 5; setup up our factors, and the twist is that we have to divide off the front number since we multiplied it in to begin with; we have to "undo" it. (x 6/3) (x 1/3) ; and the larger "factor" gets the negative (x - 6/3) (x +1/3) ; and reduce (x - 2) (x +1/3) ; we can stop here to find the zeroes , do you see why?

OpenStudy (amistre64):

step 4 can be generalized in it all, but when the first number is a "1" its kinda pointless; 1(last#) = last#

OpenStudy (anonymous):

oh i see... :)

OpenStudy (anonymous):

wait actually im confused... now i need to solve for zero right? But how do I do that?

OpenStudy (amistre64):

if we want to continue thru to get it "factored" we would go one step further; step 6 :) 2 (x - 2) (x +1/3) ; recall the the denominator came from the front of the equation to begin with, so if there is a denom left after reducing, stick it back in front for a factored form 2 (x - 2) (x +1/3) 2 (x - 2) (3x +1) <--- see what i did?

OpenStudy (amistre64):

how do you solve for zero when you are multiplying 2 numbers together? when either number is a zero, the product equals 0 right?

OpenStudy (anonymous):

yes... so how do u get a number answer?

OpenStudy (amistre64):

(x - 2) * (x +1/3) = 0 =0? =0?

OpenStudy (amistre64):

equate both "factors" to equal zero to determine when one or the other equlas zero this is simply turning it into a "multiplying by zero" effect

OpenStudy (amistre64):

5(0) = 0 a(0) = 0 ♫(0) = 0

OpenStudy (amistre64):

when does x-2 = 0? when does x+ 1/3 = 0?

OpenStudy (anonymous):

so its x-2=0 -> x=2 and x+1/3=0> x=-1/3 ???

OpenStudy (amistre64):

yep; see.. let x = 2 then: (x - 2) * (x +1/3) (2 - 2) * (2 +1/3) (0) * (2 +1/3) = 0 right? let x = -1/3 then: (-1/3 - 2) * (-1/3 +1/3) (-1/3 - 2) * (0) = 0 right?

OpenStudy (anonymous):

ok... that's what i did... YAY! Thank you!!! I really appreciate you helping me!

OpenStudy (amistre64):

youre welcome, good luck :)

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

i got questions!

OpenStudy (anonymous):

use the distributive property to simplify the expression 7(-5x+7)

OpenStudy (anonymous):

You should post questions on the math feed from now on... but the anser is... -35x-7 :)